# SecureBERT 2.0: Advanced Language Model for Cybersecurity Intelligence

### v1

@@ 1 Introduction
Domain-specific models such as the original SecureBERT have already demonstrated the value of adapting language models to cybersecurity. However, the threat landscape has evolved: modern security workflows increasingly require a deeper understanding of technical text, structured documentation, and raw code, often in combination. This shift calls for a new generation of models that can capture richer context, longer-range dependencies, and subtle relationships specific to security data.
@@ 4.2 Named Entity Recognition (NER)
Size Approximately 3.4k rows in the combined train and validation sets, and 717 rows in the test set.
@@ 5.1 Document Embedding Model
Document embedding represents a core capability for numerous cybersecurity applications that depend on semantic similarity search. By effectively capturing the underlying meaning of textual data, high-quality embeddings facilitate accurate information retrieval, classification, and threat analysis across a variety of cybersecurity tasks. We assessed SecureBERT 2.0’s ability to produce meaningful and discriminative embeddings using both bi-encoder and cross-encoder architectures. Its performance was benchmarked against two strong baselines: AttackBERT[2], a domain-adapted transformer designed for cybersecurity text, and all-MiniLM-L6-v2[3], a widely adopted, efficient sentence embedding model.
Performance was assessed using Recall at 1 (R@1), which measures the proportion of queries for which the top-ranked retrieved document is relevant, and mean Average Precision (mAP), which captures ranking quality across all retrieved results.
@@ 5.1.1 Cross-Encoder Results
The cross-encoder architecture allows for deep joint encoding of the query and candidate document, enabling nuanced semantic matching. As shown in Table 7, our model achieves an R@1 of 88.72% and an mAP of 92.14%, outperforming both baselines by a large margin. AttackBERT and all-MiniLM-L6-v2, while competitive, trail by over 10 percentage points in both metrics. This demonstrates that SecureBERT 2.0’s specialized architecture and pretraining result in a far superior understanding of cybersecurity text.
@@
Model Type R@1 (%) mAP (%)
AttackBERT 74.39 80.40
all-MiniLM-L6-v2 77.87 82.66
SecureBERT 2.0 (Cross-Encoder) 88.72 92.14
@@ 5.1.2 Bi-Encoder Results
The bi-encoder (dual encoder) configuration encodes queries and documents independently, offering much greater computational efficiency—crucial for real-time and large-scale retrieval in RAG applications. Table 8 shows that SecureBERT 2.0 not only retains but improves its retrieval performance in this more scalable setup, achieving an R@1 of 88.83% and an mAP of 92.39%. This result is particularly notable, as dual encoder models typically trade off some accuracy for speed, but SecureBERT 2.0’s robust embeddings allow it to excel in both respects. Again, both AttackBERT and all-MiniLM-L6-v2 fall significantly short.
@@
Model Type R@1 (%) mAP (%)
AttackBERT 74.39 80.40
all-MiniLM-L6-v2 77.87 82.66
SecureBERT 2.0 (Bi-Encoder) 88.83 92.39
@@ 5.1.3 Analysis and Discussion
SecureBERT 2.0’s consistent outperformance across both encoder architectures highlights several strengths:
  • Domain Adaptation: The significant margin over AttackBERT[2] and all-MiniLM-L6-v2 underscores the importance of domain-adapted pretraining. SecureBERT 2.0’s exposure to a massive, diverse cybersecurity corpus enables it to capture subtle terminology, relationships, and context unique to this field.
  • Robustness Across Architectures: Achieving comparable results with both cross- and bi-encoder architectures is rare, as the latter is typically less expressive. SecureBERT 2.0’s high-quality embeddings allow it to excel even when query and document are encoded independently, which is essential for real-world deployment at scale.
  • Implications for RAG and Real-Time Search: The bi-encoder’s performance and efficiency make SecureBERT 2.0 especially well-suited for retrieval-augmented generation, continuous threat monitoring, and interactive analyst workflows where latency and throughput are critical.
  • Semantic Precision: The high mAP values indicate that SecureBERT 2.0 not only retrieves the correct top document but also ranks relevant documents highly throughout the candidate list, providing more reliable and comprehensive results to downstream systems.

### v2

@@ 1 Introduction
Domain-specific models such as the original SecureBERT [1] have already demonstrated the value of adapting language models to cybersecurity. However, the threat landscape has evolved: modern security workflows increasingly require a deeper understanding of technical text, structured documentation, and raw code, often in combination. This shift calls for a new generation of models that can capture richer context, longer-range dependencies, and subtle relationships specific to security data.
@@ 4.2 Named Entity Recognition (NER)
Size Contains 3.4k samples in the combined train set and 717 rows for testing.
@@ 5.1 Document Embedding Model
We evaluated SecureBERT 2.0’s ability to generate meaningful and discriminative embeddings using both bi-encoder and cross-encoder architectures. Performance was benchmarked against several strong baselines, including AttackBERT[3], all-MiniLM-L6-v2[4], and other high-performing models from the Hugging Face model zoo.
Evaluation metrics include mean Average Precision (mAP), Recall at 1 (R@1), Mean Reciprocal Rank at 10 (MRR@10), and Normalized Discounted Cumulative Gain at 10 (NDCG@10), providing a comprehensive assessment of retrieval and ranking quality.
@@ 5.1.1 Cross-Encoder Results
The cross-encoder jointly encodes the query and document, allowing fine-grained interaction and semantic matching. As shown in Table 7, SecureBERT 2.0 achieves a remarkable mAP of 0.955 and R@1 of 0.948, outperforming all baselines across every metric. In particular, it surpasses the well-established ms-marco-MiniLM-L6-v2 and TinyBERT-L2-v2 models by over 3–5 percentage points on mAP and R@1. The improvements in NDCG@10 (0.986) and MRR@10 (0.983) further highlight its superior ranking and semantic precision, demonstrating exceptional understanding of cybersecurity-related text.
@@
Model mAP R@1 NDCG@10 MRR@10
BAAI/bge-reranker-large 0.741 0.741 0.894 0.864
bge-reranker-base 0.781 0.709 0.896 0.868
cross-encoder/ms-marco-MiniLM-L6 0.882 0.859 0.950 0.937
cross-encoder/ms-marco-MiniLM-L4-v2 0.904 0.863 0.964 0.956
cross-encoder/ms-marco-TinyBERT-L2-v2 0.920 0.849 0.964 0.955
SecureBERT 2.0 0.955 0.948 0.986 0.983
@@ 5.1.2 Bi-Encoder Results
The bi-encoder configuration encodes queries and documents independently, providing higher scalability and lower inference latency—ideal for large-scale retrieval and real-time RAG pipelines. Table 8 shows that SecureBERT 2.0 attains an exceptional mAP of 0.951 and R@1 of 0.984, outperforming all baselines, including domain-specific and large transformer models. Notably, despite the typical expressivity gap between bi- and cross-encoders, SecureBERT 2.0 closes this gap almost entirely, achieving near-identical ranking performance while being significantly faster in retrieval operations.
@@
Model mAP R@1 MRR@10
krlng/sts-GBERT-bi-encoder 0.672 0.806 0.852
sentence-transformers/all-MiniLM-L6-v2 0.898 0.917 0.940
basel/ATTACK-BERT 0.906 0.918 0.944
sentence-transformers/all-MiniLM-L12-v2 0.912 0.924 0.945
SecureBERT 2.0 0.951 0.984 0.989
@@ 5.1.3 Analysis and Discussion
SecureBERT 2.0 demonstrates state-of-the-art retrieval and ranking performance across both encoder architectures. Several key observations emerge:
  • Superior Domain Adaptation: The model’s advantage over AttackBERT and MiniLM variants reinforces the benefits of pretraining on a large-scale, diverse cybersecurity corpus. SecureBERT 2.0 captures nuanced terminology and contextual dependencies unique to the cybersecurity domain.
  • Cross-Architecture Consistency: Achieving near-parity between cross- and bi-encoder results is rare, underscoring the embedding robustness and semantic depth of SecureBERT 2.0. This flexibility enables both accuracy-sensitive and latency-sensitive deployment scenarios.
  • Scalability and Efficiency: The bi-encoder’s superior R@1 and mAP make SecureBERT 2.0 ideal for high-throughput applications such as retrieval-augmented generation (RAG), automated threat intelligence systems, and analyst-support tools where fast, accurate retrieval is essential.
  • Ranking Precision: The consistently high MRR and NDCG values indicate that SecureBERT 2.0 not only retrieves relevant results but also ranks them effectively, ensuring downstream models and analysts receive the most contextually relevant information first.